• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 00:04:55 on October 23, 2005

    Entry number 157735

    keyword=Shift summary

    Shift Crew:  Sawatzky, Hansen


    Comments:

    Shift starts with MCC requesting a little time to send pulsed beam into hall to run some diagnostics.

    At about 16:30 Q2 and Q3 in the right arm tripped. Q1 tripped while I was down in the hall resetting Q3. Q3 refused to ramp back up so Tech (Ed) was called. He arrived with a plan and valve controller card in hand and Jack arrived to help with the install.

    Once the magnets stabilized (around 19:30) it was pretty smooth through the rest of the shift.


    Events:
    16:00 MCC still working on beam quality -- a few pairs taken while MCC works.

    16:10 MCC asked to take CW beam away briefly to run pulsed beam into our Hall to run some tests.

    16:30 Right Q2/Q3 tripped, followed by Q1 5-10 minutes later while I was in the hall resetting Q3 (controlled access). All quads were being very flaky. Q3 never did want to ramp so tech (Ed) was called in. He believes the problem is due to a bad valve controller that was suspect earlier href="http://www.jlab.org/~adaq/halog/html/0510_archive/051022120605.html">157681. He arrived with a replacement in hand.

    16:35 MCC calls to tell us that the beam looks OK and asks if they can ramp up the current. Of course we're now trying to reset our magnets. Argh.

    18:40 Magnets are up. Ed asks us to cycle Q2/Q3 while he stays in the hall. See· 157710. Thanks to Ed and Jack for coming in!

    19:35 Q2 tripped two more times when attempting to cycle it (once to 1600A, again to 1500A). These trips require access to the hall to reset. Decided to NOT cycle Q2 -- we just ramped it directly to its production setpoint of ~1350A. We're (finally) ready for beam.

    19:50 Requested 10 uA, did a spot++ run and had MCC work on centering the beam.

    20:00 Started first good production run. Everything looks to be OK.

    21:20 Beam jumped 2mm on BPMB-Y we caught the problem and had MCC correct it at 20:30 (Run 5819).

    22:35 MCC takes beam away for their PZT test.

    22:55 Beam is back.

    23:10 Noticed that Compton coincidence rate ('Both') was stuck at zero. Restarted Compton CODA session and began new run. Pair rates back to ~100Hz.


    Run Summary:
    Beam Energy: 3.1785 GeV (Set Value), 3.1816 GeV (Meas)
    Target: 20cm Hydrogen
      Raster: On (4mm x 2mm)
      Left HRS Setting: 6 deg, 3.1234 GeV/c, septum set point: 323.61 A
      Right HRS Setting: 6 deg, 3.1l26 GeV/c, septum set point: 316.92A


    Note: Q2 on Right arm was not cycled to 1600A after its last trip. See comments.
    5818  Production: 41.2k pairs
    5819  Production: 36.3k pairs, 2mm beam shift on BPMB-Y for 10 minutes before we caught it...
    5820  Production: 32.2k pairs, stopped a little early for PZT test
    5821  Production: XXX pairs
    5822  Production: hand-off to next shift